OpenStudy (anonymous):

which ordered pair is a solution of the equation y=x-2 (3,1) (1,3) (-1,3) (3,-1)

pooja195 (pooja195):

Plug in your opitons.

OpenStudy (anonymous):

example ? @pooja195

pooja195 (pooja195):

A. 1=(3)-2 B. 3=(1)-2 C. 3=(-1)-2 D.-1=(3)-2 See which one is true.

OpenStudy (anonymous):

1=(3)-2 ? @pooja195

pooja195 (pooja195):

Yep

OpenStudy (anonymous):

So it would be (1,3) @pooja195

pooja195 (pooja195):

Yes

OpenStudy (anonymous):

Can you help me with another one ? @pooja195

OpenStudy (anonymous):

Which ordered pair is a solution of the equation y=5x? (-2,10) (-5,25) (-3,15) (-2,-10)

pooja195 (pooja195):

Same thing

OpenStudy (anonymous):

how do I plug in my choices @pooja195

pooja195 (pooja195):

(x,y) the left number is x right is y

OpenStudy (anonymous):

I don't get it ??? @pooja195